OpenStudy (anonymous):

A metal conduit will be used as a pathway for wiring through a concrete block. The conduit is a 6 foot long rod with an outer diameter of 2.8 inches and an inner diameter of 2.6 inches. What is the volume of metal in the conduit in cubic centimeters?

OpenStudy (mathstudent55):

Do you know how to find the volume of a cylinder.

OpenStudy (anonymous):

2*pi*r

OpenStudy (mathstudent55):

That is the circumference of a circle. You need the volume of a cylinder.

OpenStudy (anonymous):

oh pi*r^2*h

OpenStudy (mathstudent55):

Great. All you need to do is find the volume of the outer dimension of the pipe, then find the volume of the inner dimension of the pipe and subtract.

OpenStudy (mathstudent55):



OpenStudy (anonymous):

61.07 is my answer?

OpenStudy (mathstudent55):

The outer cylinder has volume: \(V = \pi R^2h = \pi \left(\dfrac{D}{2}\right)^2h = \dfrac{\pi D^2h}{4}\) The inner cylinder has volume: \(V = \pi r^2h = \pi \left(\dfrac{d}{2}\right)^2h = \dfrac{\pi d^2h}{4}\) The voluyme of the hollow pipe is: \(V = \dfrac{\pi D^2h}{4} - \dfrac{\pi d^2h}{4}\) \(V = \dfrac{\pi (D^2 - d^2)h}{4}\)

OpenStudy (mathstudent55):

You are correct that the volume is 61.07 cu in. The question asks for the volume in \(cm^3\).

OpenStudy (anonymous):

ok i got it now thank you so much

OpenStudy (mathstudent55):

wlcm What do you get in \(cm^3\)?

OpenStudy (anonymous):

1001 cm^3

OpenStudy (mathstudent55):

Excellent!! Good job.